OpenStudy (sagewilson):

Multiply or divide as indicated. Simplify your answer, if possible. (Do not express answer in exponential form.) Expression below.

OpenStudy (sagewilson):

\[\frac{ 10 }{ 10^{-2}}\] I think the answer would be \[10^{-2}\] but the answer has to be 4 numbers long.

OpenStudy (sagewilson):

the - and the 2 are one thing so the answer 10^-2 is 3 spaces, not 4.

OpenStudy (sapphiremoon):

What you have is 10 divided by 10^-2. 10^-2 is actually 1/10^2, or 1/100. So dividing 10 by 1/100 requires some fractional division if your scientific calculator doesn't do that for you: \[\frac{ 10 }{ 1 } \times \frac{ 100 }{ 1 } = 1000\] Because with fractional division you multiply by the reciprocal of the second fraction. Make sense?

OpenStudy (welshfella):

use the laws of exponents it equal 10^1 / 10^-2 when you divide as above what do you do with the exponents?

OpenStudy (welshfella):

once you have the answer in exponent form convert to normal numbers

OpenStudy (sagewilson):

@SapphireMoon Yes I understand now. @welshfella Well in the lesson I'm given if you have something like x^3 / x^-2 I'd end up getting x^1.

OpenStudy (welshfella):

x^3 / x^-2 = x^(3-(-2) = x^5

OpenStudy (welshfella):

10^1 / 10^-2 = 10^3 = 1000

OpenStudy (sagewilson):

I was taught to add the exponents not multiply them. I just looked back at the lesson. That was the rule for the addition and subtraction one. I was using the wrong equation.

OpenStudy (welshfella):

you add the exponents when you are multiplying the exponential terms when dividing the terms you subtract the exponents.

OpenStudy (sapphiremoon):

You don't multiply them unless it's something like: \[(10^{2})^{3} = 10^{6}\] You add for multiplication of exponents and subtract for division of exponents. So 1 - (-2) = 3 because subtracting a negative number means adding a positive one.

OpenStudy (welshfella):

10^4 * 10 ^2 = 10^6 10^4 / 10^2 = 10^(4-2) = 10^2

OpenStudy (sagewilson):

These are some examples my lesson is giving me. The thing is if I put this it wont be enough to actually fill in the amount of letters I need for the full answer.

OpenStudy (sagewilson):

It does look like I have to add though.

OpenStudy (welshfella):

Manipulating exponential terms comes easy with practice. You just need to know the 3 main rules.

OpenStudy (welshfella):

where do you think you have to add? In the question you posted?

OpenStudy (sagewilson):

Looking back at the lesson, it looks like I have to subtract instead of add since the numerators exponent is positive and the denominators is negative.

OpenStudy (sagewilson):

Since it's a division problem. If it were a multiplication problem I'd add.

OpenStudy (welshfella):

Yes When its a multiplication problem you add exponents division - you subtract exponents.

OpenStudy (welshfella):

the fact that the numerator is positive and denominator is a negative has nothing to do with the choice.

OpenStudy (sapphiremoon):

^^

OpenStudy (sagewilson):

Okay, I know that now. So the answer would be 1,000 like @SapphireMoon said since I'm solving it.

OpenStudy (sapphiremoon):

Yep!

OpenStudy (welshfella):

yes I think they want you to use exponent processing first then convert to normal form later. But converting to normal form first then simplifying gives the same answer.

OpenStudy (welshfella):

- but i may be wrong. Ill think you'll get the same credit whichever way.

OpenStudy (sagewilson):

Thanks to both of you for all of your help.

OpenStudy (sapphiremoon):

You're welcome! I remember struggling with this in prealgebra (or whenever it was, a very long way from precal anyways), happy to help!

OpenStudy (welshfella):

yw

OpenStudy (welshfella):

If few more facts about exponents x can be written as x^1 6 = 6^1 abd so on. x^0 ( in fact any variable or number with exponent 0) = 1

OpenStudy (welshfella):

y^0 = 1 , 1563^0 = 1 etc